OpenStudy (anonymous):

Use algebra to find the inverse of the given one-to-one function. Can someone show me how to solve this problem? f(x) = (x^5 + 9)^3

OpenStudy (anonymous):

just like if you were solving \[(x^5+9)^3=12\] for x, solve \[(y^5+9)^3=x\] for y

OpenStudy (anonymous):

\[(y^5+9)^3=x\] \[y^5+9=\sqrt[3]{x}\] \[y^5=\sqrt[3]{x}-9\] \[y=\sqrt[5]{\sqrt[3]{x}-9}\]

OpenStudy (anonymous):

Thank you!